OpenStudy (anonymous):

inverse of the function f(x)=3x+9? would this be : f(x)=x-9/3 ?

jimthompson5910 (jim_thompson5910):

do you mean \[\Large x - \frac{9}{3}\] OR \[\Large \frac{x-9}{3}\] ??

OpenStudy (anonymous):

the second one

jimthompson5910 (jim_thompson5910):

you have the correct inverse function, you would write it like this \[\Large f^{-1}(x) = \frac{x-9}{3}\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome